### v1

# Inhomogeneous Kinetic Equation for Mixed Neutrinos: Tracing the Missing Energy

Abstract
Flavor-dependent neutrino transport is described by a well-known kinetic equation for occupation-number matrices in flavor space. However, as an overlooked theoretical problem, we show that in the inhomogeneous case, neutrino-neutrino refractive energy is not conserved. We derive the missing gradient terms in the fast flavor limit (vanishing neutrino masses), and prove that the missing refractive energy is traded with the huge reservoir of neutrino kinetic energy through gradients of the weak interaction potential. Even small changes of the kinetic energy accommodate the refractive energy gained or lost. Flavor evolution alone is negligibly affected by the new terms.
Introduction.—Flavor evolution in a dense neutrino environment is often described by a kinetic equation for the flavor density matrices $\rho({\bf p},{\bf x},t)$, where the diagonal entries are the usual occupation numbers Dolgov:1980cq ; Rudsky ; Sigl:1993ctk ; Sirera:1998ia ; Yamada:2000za ; Vlasenko:2013fja ; Volpe:2013uxl ; Serreau:2014cfa ; Kartavtsev:2015eva ; Fiorillo:2024wej . In the homogeneous case, they are defined as the expectation values of the number-operator matrices ${\cal D}_{\alpha\beta}=a^{\dagger}_{\beta,{\bf p}}a_{\alpha,{\bf p}}$ with the flavor indices $\alpha$ and $\beta$ and $a^{\dagger}_{\alpha,{\bf p}}$ the creation operator for a left-handed neutrino of flavor $\alpha$ with momentum ${\bf p}$. In the limit of vanishing neutrino masses and flavor mixing, nontrivial flavor evolution can still arise through classical instabilities that would engender strong flavor correlations Sawyer:2015dsa ; Chakraborty:2016lct ; Izaguirre:2016gsx ; Tamborra:2020cul ; Richers:2022zug . Dubbed fast flavor evolution, this subject has been intensely studied in view of practical consequences for neutrino flavor transport in stellar core collapse and binary neutron star mergers Xiong:2022vsy ; Fernandez:2022yyv ; Ehring:2023abs ; Ehring:2023lcd ; Nagakura:2023mhr ; Nagakura:2023xhc ; Zaizen:2023wht ; Grohs:2023pgq ; Abbar:2023zkm ; Akaho:2023brj ; Froustey:2023skf ; Nagakura:2023wbf .
Instead, we turn to an overlooked theoretical problem of the usual equations of motion (EOMs): their lack of energy conservation. In their often-used form, they are
$\bigl{(}\partial_{t}+{\bf v}\cdot{\bm{\nabla}}_{\bf r}\bigr{)}\rho_{{\bf p},{\bf r},t}=-i\bigl{[}{\sf H}_{{\bf p},{\bf r},t},\rho_{{\bf p},{\bf r},t}\bigr{]},$ (1)
where ${\sf H}_{{\bf p},{\bf r},t}=\sqrt{2}G_{\rm F}\sum_{{\bf p}^{\prime}}\rho_{{\bf p}^{\prime},{\bf r},t}\,(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})$ is a matrix that drives flavor evolution by neutrino-neutrino refraction. The so-called advection term proportional to ${\bf v}={\bf p}/|{\bf p}|$ quantifies a drift in coordinate space caused by inhomogeneities. Among other consequences, this term allows for ${\bf p}$-dependent instabilities and even an initially homogeneous system can develop flavor disturbances that grow exponentially and drift to ever smaller scales Duan:2014gfa ; Mirizzi:2015fva ; Bhattacharyya:2020jpj ; Johns:2020qsk ; Tamborra:2020cul ; Richers:2021nbx ; Bhattacharyya:2022eed .
As detailed later, even an elementary example of two colliding beams consisting initially of $\nu_{e}$ and $\nu_{x}$ reveals that the refractive interaction energy changes dramatically as soon as initial seeds of inhomogeneity grow. The solution of this puzzle is that one cannot look at flavor evolution in isolation. Inhomogeneities in the weak potential exert a force and trade refractive interaction energy with the huge reservoir of neutrino kinetic energy. We here derive the missing terms in Eq. (1) and show that to lowest order in a gradient expansion, the energy exchange can be exactly accounted for.
Interaction energy.—We first need to define the $\nu$$\nu$ interaction energy. The starting point is the many-body Hamiltonian $\mathcal{H}=\mathcal{H}_{0}+\mathcal{U}$, where $\mathcal{H}_{0}=\sum_{\alpha,{\bf p}}\epsilon_{\bf p}a^{\dagger}_{\alpha,{\bf p}}a_{\alpha,{\bf p}}$ is the kinetic energy operator with $\epsilon_{\bf p}=|{\bf p}|$ in the massless limit. Moreover, the interaction energy is
$\mathcal{U}$ $=$ $\frac{\sqrt{2}G_{\rm F}}{8}\sum_{\left\{{\bf p}\right\},\alpha,\beta}a^{\dagger}_{\alpha,{\bf p}_{1}}a_{\alpha,{\bf p}_{2}}a^{\dagger}_{\beta,{\bf p}_{3}}a_{\beta,{\bf p}_{4}}$ (2)
×$\kern 70.0001pt{}\times\overline{u}_{{\bf p}_{1}}\gamma^{\mu}u_{{\bf p}_{2}}\overline{u}_{{\bf p}_{3}}\gamma_{\mu}u_{{\bf p}_{4}},$
where $G_{\rm F}$ is Fermi’s constant, $\sum_{\left\{{\bf p}\right\}}$ is performed over all momenta such that ${\bf p}_{1}+{\bf p}_{3}={\bf p}_{2}+{\bf p}_{4}$, and $u_{\bf p}$ is the spinor of a left-handed massless neutrino with momentum ${\bf p}$, normalized such that $\overline{u}_{\bf p}\gamma^{0}u_{\bf p}=2$.
The energy of the system is the expectation value of $\mathcal{H}$ over the quantum state. In particular, the kinetic energy is $K=\langle\mathcal{H}_{0}\rangle=\int d^{3}{\bf r}\sum_{\bf p}\epsilon_{\bf p}\mathrm{Tr}(\rho_{\bf p})$. For the interaction energy, we need to determine the average value of a string of four creation and annihilation operators. We do so in the mean field, or Hartree-Fock, approximation, where the quantum state is assumed to be the product of single-particle states, so that the expectation value factorizes as $\langle a^{\dagger}_{1}a_{2}a^{\dagger}_{3}a_{4}\rangle=\langle a^{\dagger}_{1}a_{2}\rangle\langle a^{\dagger}_{3}a_{4}\rangle+(\delta_{2,3}-\langle a^{\dagger}_{3}a_{2}\rangle)\langle a^{\dagger}_{1}a_{4}\rangle$. In this way we define a potential energy of interaction $U=\langle\mathcal{U}\rangle$. For a homogeneous system, it is
$U$ $=$ $\frac{\sqrt{2}G_{\rm F}}{2}\int d^{3}{\bf r}\sum_{{\bf p},{\bf p}^{\prime}}\Bigl{[}\mathrm{Tr}(\rho_{\bf p})\,\mathrm{Tr}(\rho_{{\bf p}^{\prime}})+\mathrm{Tr}(\rho_{\bf p}\rho_{{\bf p}^{\prime}})\Bigr{]}$ (3)
×$\kern 80.00012pt{}\times(1-\cos\theta_{{\bf p},{\bf p}^{\prime}}),$
where we neglect a single-particle term constituting an irrelevant (infinite) renormalization of the neutrino chemical potential.
In the two-flavor case, the density matrices are often decomposed into their trace and trace-free part as
$\rho_{\bf p}=\frac{P_{\bf p}^{0}\,\mathbb{1}+{\vec{P}}_{\bf p}\cdot{\vec{\sigma}}}{2},$ (4)
where $\mathbb{1}$ is a unit matrix, $P_{\bf p}^{0}=\mathrm{Tr}(\rho_{{\bf p}})$, $\vec{\sigma}$ a vector of Pauli matrices, and $\vec{P}_{\bf p}$ a polarization vector. ($\vec{P}$ is a vector in flavor space, ${\bf p}$ one in phase space.) Exactly one particle in mode ${\bf p}$ implies $P_{\bf p}^{0}=|\vec{P}_{\bf p}|=1$. Only $\vec{P}_{\bf p}$ evolves by oscillations, whereas the trace (total particle number in a ${\bf p}$-mode) is conserved. The oscillatory part of the interaction energy is
$U_{\mathrm{osc}}=\frac{\sqrt{2}G_{\rm F}}{4}\int d^{3}{\bf r}\sum_{{\bf p},{\bf p}^{\prime}}(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})\,{\vec{P}}_{\bf p}\cdot{\vec{P}}_{{\bf p}^{\prime}},$ (5)
corresponding to that part of Eq. (3) that depends on the trace-free parts of the $\rho$ matrices.
For simplicity we assume in the following a system that remains homogeneous in all but the spatial $z$-direction and use $v=v_{z}$. The standard EOM Eq. (1) falls into pieces for particle number and flavor polarization as
$\partial_{t}P^{0}_{\bf p}+\,v\partial_{z}P^{0}_{\bf p}$ $=$ $0,$ (6a)
$\partial_{t}\vec{P}_{\bf p}+v\partial_{z}\vec{P}_{\bf p}$ $=$ $\sqrt{2}G_{\rm F}\bigl{(}{\vec{P}}_{0}-v{\vec{P}}_{1}\bigr{)}\times{\vec{P}}_{\bf p},$ (6b)
where we use the angular moments ${\vec{P}}_{0}=\sum_{\bf p}{\vec{P}}_{\bf p}$ and ${\vec{P}}_{1}=\sum_{\bf p}v{\vec{P}}_{\bf p}$. Thus, the usual equations decouple the neutrino number from their flavor evolution.
Different from a neutrino gas subject to collisions with matter (see, e.g., Refs. Johns:2021qby ; Nagakura:2022qko ; Xiong:2022zqz ; Fiorillo:2024 ; Liu:2023pjw ; Lin:2022dek ; Johns:2022yqy ; Shalgar:2022rjj ; Padilla-Gay:2022wck ; Fiorillo:2023ajs ; Johns:2023xae ), where energy is not conserved anyway, Eqs. (Inhomogeneous Kinetic Equation for Mixed Neutrinos: Tracing the Missing Energy) describe a closed system that admits a conserved energy. The kinetic part $K=\int d^{3}{\bf r}\sum_{\bf p}\epsilon_{\bf p}P^{0}_{\bf p}$ is seen to be conserved from Eq. (6a). Also the nonoscillatory part of the interaction energy is separately conserved. $U_{\mathrm{osc}}=(\sqrt{2}G_{\rm F}/4)\int d^{3}{\bf r}\,({\vec{P}}_{0}^{2}-{\vec{P}}_{1}^{2})$, which follows from Eq. (5), must then be conserved as well. In the homogeneous case, this is indeed the case; the dynamics is periodic and described by a fictitious pendulum whose conserved energy coincides with $U_{\mathrm{osc}}$ Johns:2019izj ; Padilla-Gay:2021haz ; Fiorillo:2023mze ; Fiorillo:2023hlk . However, Eq. (6b) immediately reveals that the conservation breaks down for inhomogeneous settings
$\frac{dU_{\mathrm{osc}}}{dt}=-\frac{\sqrt{2}G_{\rm F}}{2}\int d^{3}{\bf r}\,\Bigl{(}{\vec{P}}_{0}\cdot\partial_{z}{\vec{P}}_{1}-{\vec{P}}_{1}\cdot\partial_{z}{\vec{P}}_{2}\Bigr{)},$ (7)
where ${\vec{P}}_{2}=\sum_{\bf p}v^{2}{\vec{P}}_{\bf p}$. Therefore, the traditional EOMs are not consistent.
Two-beam example.—The nonconservation of $U_{\mathrm{osc}}$ is a large effect as we show in a simple example of two opposite beams along the $z$-axis ($v=\pm 1$). The energies and number densities are equal, one beam initially occupied with $\nu_{e}$, the other with $\nu_{x}$. As usual, the beams represent many neutrinos with ${\bf p}$’s close enough that the small spread is unimportant on the relevant time scales. Therefore, we represent the flavor polarization of each beam by a highly occupied $\rho$ matrix and concomitant polarization vector ${\vec{P}}_{\pm}(z,t)$. We thus need to solve
$(\partial_{t}+\partial_{z}){\vec{P}}_{+}$ $=$ $2\,{\vec{P}}_{-}\times{\vec{P}}_{+},$ (8a)
$(\partial_{t}-\partial_{z}){\vec{P}}_{-}$ $=$ $2\,{\vec{P}}_{+}\times{\vec{P}}_{-},$ (8b)
where the interaction strength was absorbed in the units of time and space. For $\partial_{z}{\vec{P}}_{\pm}=0$ there is no instability, but the system has unstable inhomogeneous modes.
We can show energy nonconservation analytically in the initial linear regime. In this limit, it is convenient to express the $x$$y$ part of the polarization vector (the off-diagonal element of the density matrix) in the form $\psi_{\pm}=P^{x}_{\pm}+iP^{y}_{\pm}$ and initially we take $P_{\pm}^{z}(z,0)=\zeta_{\pm}$ to be constant. The linearized EOMs ($|\psi|\ll|\zeta|$) are
$\bigl{(}\partial_{t}+\partial_{z}\bigr{)}\,\psi_{+}$ $=$ $2\,i\bigl{(}+\zeta_{-}\psi_{+}-\zeta_{+}\psi_{-}\bigr{)},$ (9a)
$\bigl{(}\partial_{t}-\partial_{z}\bigl{)}\,\psi_{-}$ $=$ $2\,i\bigl{(}-\zeta_{-}\psi_{+}+\zeta_{+}\psi_{-}\bigr{)}.$ (9b)
[…]
$\kern-20.00003pt\begin{pmatrix}\widetilde{\psi}_{+,k}(t)\\ \widetilde{\psi}_{-,k}(t)\end{pmatrix}$ $=$ $\alpha_{k}\begin{pmatrix}k+\omega_{0}+i\gamma_{k}\\ k-\omega_{0}-i\gamma_{k}\end{pmatrix}e^{(\gamma_{k}-i\omega_{0})t}$ (17)
${}+\beta_{k}\begin{pmatrix}k+\omega_{0}-i\gamma_{k}\\ k-\omega_{0}+i\gamma_{k}\end{pmatrix}e^{-(\gamma_{k}+i\omega_{0})t}.$
Matching this expression with the initial conditions $\widetilde{\psi}_{\pm,k}(0)=\widetilde{\psi}_{\pm,k}^{0}$ reveals for the growing piece
$\alpha_{k}=\frac{(\gamma_{k}+i\omega_{0})\left(\widetilde{\psi}_{+,k}^{0}+\widetilde{\psi}_{-,k}^{0}\right)-ik\left(\widetilde{\psi}_{+,k}^{0}-\widetilde{\psi}_{-,k}^{0}\right)}{4k\gamma_{k}}.$ (18)
Substituting in the transverse part of the oscillation energy, we finally find
$\Delta U_{\mathrm{osc}}(t)=L\sum_{k=k_{1}}^{k_{2}}2k(k-\overline{k})|\alpha_{k}|^{2}e^{2\gamma_{k}t},$ (19)
where we restrict the summation to the range of unstable eigenmodes. Notice that the most unstable eigenmode $k=\overline{k}$ does not contribute to energy nonconservation due to the vanishing prefactor. Otherwise, an exponential growth appears already at the linear level.
For more than two beams, from Eq. (7) it follows that $dU_{\mathrm{osc}}/dt\propto\sum_{k}kv{\vec{P}}_{v}(k)\cdot{\vec{P}}_{v^{\prime}}^{*}(k)(1-vv^{\prime})$, and thus it is still true that the energy change grows with $e^{2\overline{\gamma}t}$, provided that $k\neq 0$; homogeneous modes, with $k=0$, do not lead to energy change due to the prefactor $kv$.
For a numerical solution that extends to the nonlinear regime, we use periodic boundary conditions on a box of length $L=100$, divided in $N$ spatial bins. As initial conditions, we use $P^{z}_{\pm}(0)=\zeta_{\pm}=\pm 1/2$. These parameters imply that the range of unstable modes is delimited by $k_{1}=-2$ and $k_{1}=0$, the maximum growth rate obtains for $\overline{k}=-1$ and is $\overline{\gamma}=1$, and the initial oscillation energy is $U_{\mathrm{osc}}(0)=\zeta_{-}\zeta_{+}L=-25$. The transverse components $P^{x,y}_{\pm}$ are seeded with randomly sampled functions with a spatial dependence
$P^{x,y}_{\pm}=\sum_{n=-N_{\mathrm{max}}}^{N_{\mathrm{max}}}c^{x,y}_{\pm,n}e^{i\phi^{x,y}_{\pm,n}+i\frac{2\pi nz}{L}},$ (20)
where the amplitudes are sampled from a normal distribution with a variance $\sigma^{2}=10^{-8}$ and the phases are uniformly sampled from 00 to $2\pi$; the functions are real and so $c_{n}=c_{-n}$ and $\phi_{n}=-\phi_{-n}$. We use arbitrarily $N_{\mathrm{max}}=100$ to avoid seeds at too small scales. Numerical stability requires keeping fluctuations at the scale of the grid spacing as small as possible.
Figure 1 shows the solution for a single realization of initial seeds. The upper panel shows contours of the $\nu_{x}$-content of the beam initially occupied with $\nu_{e}$, as a function of $z$ and $t$. The lower panel shows $U_{\mathrm{osc}}(t)$ in units of $U_{\mathrm{osc}}(0)$. On the shown timescale, the range of unstable modes $-2<k<0$ grows nonlinear and then keeps oscillating, with beats between different modes causing an irregular pattern that in detail depends on the choice of seeds. In the longer run, these nonlinear modes feed higher-$k$ modes and flavor variations will obtain on ever smaller scales. Either way, the nonconservation of $U_{\mathrm{osc}}$ is a large effect, not a small correction.
Figure 1: Solution of our two-beam example for one realization of initial seeds. Top: Contours of $\nu_{x}$ content of $v=+1$ beam as a function of $z$ and $t$. Bottom: Evolution of oscillation energy $U_{\mathrm{osc}}(t)$ in units of $U_{\mathrm{osc}}(0)$.
[…]
To augment Eq. (1) with the missing terms, we note that in an inhomogeneous setting, the space-dependent occupation-number matrix is defined by a Wigner distribution Wigner:1932eb , the expectation value of
$\mathcal{D}_{\alpha\beta}({\bf r},{\bf p})=\sum_{\bf k}a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a_{\alpha,{\bf p}-{\bf k}/2}e^{-i{\bf k}\cdot{\bf r}}.$ (21)
Here, the Fourier wavevector ${\bf k}$ measures the typical length scale over which the density matrix is changing, and is therefore $|{\bf k}|\sim\ell^{-1}\ll|{\bf p}|$. The time derivative is obtained using Heisenberg’s equations. The commutator with the kinetic energy is easily established as
$\left(\frac{\partial\mathcal{D}_{\alpha\beta}}{\partial t}\right)_{0}$ $=$ $i\sum_{\bf k}a^{\dagger}_{\beta,{\bf p}+\frac{{\bf k}}{2}}a_{\alpha,{\bf p}-\frac{{\bf k}}{2}}e^{-i{\bf k}\cdot{\bf r}}\left(\epsilon_{{\bf p}+\frac{{\bf k}}{2}}-\epsilon_{{\bf p}-\frac{{\bf k}}{2}}\right)$ (22)
$\simeq$ $-{\bf v}\cdot{\bm{\nabla}}_{\bf r}\mathcal{D}_{\alpha\beta},$
recovering the usual advection term.
For the interaction part, the calculations are much lengthier, so we here only outline the general strategy and report details in the Supplemental Material (SM). The commutator $[\mathcal{D},\mathcal{U}]$ produces strings of four creation and annihilation operators, or four-point correlation functions. The average of these combinations are evaluated again using the mean-field approximation. In the homogeneous limit, when taking the expectation value of $\langle a^{\dagger}_{1}a_{2}\rangle$, the momenta of state $1$ and $2$ must be equal. We go one step further and include the inhomogeneity to first order in ${\bf k}$. The final result, fully derived in the SM, is
$\partial_{t}\rho_{{\bf p},{\bf r}}+{\bf v}\cdot{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}+\frac{1}{2}\left\{{\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}\right\}-\frac{1}{2}\left\{{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\right\}=i\Bigl{[}\rho_{{\bf p},{\bf r}},{\sf\Omega}_{{\bf p},{\bf r}}\Bigr{]},$ (23)
where we introduce
${\sf\Omega}_{{\bf p},{\bf r}}^{(0)}=\sqrt{2}G_{\rm F}\sum_{{\bf p}^{\prime}}(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})\Bigl{[}\rho_{{\bf p}^{\prime},{\bf r}}+\mathrm{Tr}(\rho_{{\bf p}^{\prime},{\bf r}})\mathbb{1}\Bigr{]}$ (24)
that coincides with the standard ${\sf H}_{{\bf p},{\bf r}}$ stated after Eq. (1) except for the additional trace term that drops out in commutators, but not in the anticommutators on the left-hand side. Moreover,
${\sf\Omega}_{{\bf p},{\bf r}}={\sf\Omega}_{{\bf p},{\bf r}}^{(0)}+\sqrt{2}G_{\rm F}\sum_{{\bf p}^{\prime}}\frac{p+p^{\prime}}{p^{2}p^{{}^{\prime}2}}\,\left({\bf p}\times{\bf p}^{\prime}\right)\cdot\Bigl{[}{\bm{\nabla}}_{\bf r}\rho_{{\bf p}^{\prime},{\bf r}}+\mathrm{Tr}({\bm{\nabla}}_{\bf r}\rho_{{\bf p}^{\prime},{\bf r}})\mathbb{1}\Bigr{]},$ (25)
where $p=|{\bf p}|$ and $p^{\prime}=|{\bf p}^{\prime}|$. The correction to the quasi-particle energy can be obtained directly from the interaction Hamiltonian, taking the mean-field Hamiltonian expanded to first order in the gradients.
The new terms are intuitive if we interpret ${\sf\Omega}_{{\bf p},{\bf r}}$ as the renormalization of the quasiparticle energy induced by the interaction, which is in general a matrix in flavor space, analogous to the renormalization of quasiparticle energy in Landau’s theory of a Fermi liquid Landau:1958joj . This induces a renormalization in the group velocity ${\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$ (only the zero order terms must be kept which already contain one gradient) and a weak force field $-{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$.
All additional terms are small, of the order of $r_{\mu}r_{l}$, and thus will not produce quantitatively large flavor-conversion effects. However, as a qualitatively new effect, the renormalized group velocity causes a slow spatial drift of neutrinos; the order of magnitude of the velocity is $|{\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}|\sim r_{\mu}$. Since $r_{\mu}\sim 10^{-11}$, this is completely negligible compared to the standard neutrino velocity.
The most significant impact comes from the weak force $-{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}}$, which can change the neutrino kinetic energy. Part of this force originates from gradients in the neutrino number density. Our new insight is that an additional component originates from gradients in flavor composition. To see this clearly, we rewrite the EOMs as
$\partial_{t}\rho_{{\bf p},{\bf r}}+{\bf v}\cdot{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}=i\bigl{[}\rho_{{\bf p},{\bf r}},{\sf\Omega}_{{\bf p},{\bf r}}\bigr{]}+{\bm{\nabla}}_{\bf r}\cdot{{\bm{\Phi}}}_{{\bf p},{\bf r}}-{\bm{\nabla}}_{\bf p}\cdot{{\bf F}}_{{\bf p},{\bf r}},$ (26)
where
${\bm{\Phi}}_{{\bf p},{\bf r}}=\frac{\sqrt{2}G_{\rm F}}{2}\sum_{{\bf p}^{\prime}}(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})\Bigl{[}\bigl{\{}\rho_{{\bf p}^{\prime},{\bf r}},{\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\bigr{\}}+2\,\mathrm{Tr}(\rho_{{\bf p}^{\prime},{\bf r}}){\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\Bigr{]}$ (27)
is the number flux of neutrinos passing through a surface element of coordinate space at phase-space location $\{{\bf p},{\bf r}\}$, whereas ${\bf F}_{{\bf p},{\bf r}}$ is the same in momentum space, given by the same expression with ${\bm{\nabla}}_{\bf p}\to{\bm{\nabla}}_{\bf r}$. Crucially, this term couples the trace of the density matrix with the polarization vector; taking the trace, we find
$\mathrm{Tr}({\bf F}_{{\bf p},{\bf r}})=\frac{\sqrt{2}G_{\rm F}}{2}\sum_{{\bf p}^{\prime}}(1-\cos\theta_{{\bf p},{\bf p}^{\prime}})\Bigl{[}3P^{0}_{{\bf p}^{\prime},{\bf r}}{\bm{\nabla}}_{\bf r}P^{0}_{{\bf p},{\bf r}}+{\vec{P}}_{{\bf p}^{\prime},{\bf r}}\cdot{\bm{\nabla}}_{\bf r}{\vec{P}}_{{\bf p},{\bf r}}\Bigr{]}.$ (28)
Thus, spatial gradients of the polarization vectors can feed $\nu$ kinetic energy. The order of magnitude of this term implies a net rate of energy gain or loss of order $d\epsilon_{\bf p}/dt\simeq\mu/\ell$. Over a timescale $\ell$, the energy that a $\nu$ can accumulate is of order $\mu\ll\epsilon_{\bf p}$, so we recover that this is a small effect relative to $\epsilon_{\bf p}$. It provides the missing channel by which kinetic and refractive energy can be traded. Even a small amount of energy $\mu$ lost (gained) from the large $\nu$ kinetic energy explains the change in the refractive energy, since $U_{\mathrm{osc}}/K\simeq\mu/\epsilon_{\bf p}$.
With the new kinetic equations, we can directly prove that to first order in $r_{\mu}$ and $r_{\ell}$, the total energy is indeed conserved because the kinetic energy $K$ is no longer conserved. After an integration by parts, we find
$\frac{dK}{dt}=\int d^{3}{\bf r}\sum_{\bf p}\mathrm{Tr}\left({\bf v}\cdot{\bf F}_{{\bf p},{\bf r}}\right).$ (29)
Substituting the polarization vector part of Eq. (28) in Eq. (29), we find that it precisely balances the rate of change in the oscillation energy found in Eq. (7).
Discussion.—We have shown that the usual kinetic equation for mixed neutrinos Eq. (1) is non-conservative in the inhomogeneous case. We have illustrated this point with a simple two-beam example, where the $\nu$$\nu$ interaction energy strongly changes, and we have also shown this effect analytically in the linear regime. Deriving small gradient terms beyond Eq. (1), we have shown that the refractive energy gained or lost is precisely traded with neutrino kinetic energy that usually is not followed. In our two-beam example, the monochromatic initial energy distribution develops small space-time dependent shifts that account for the missing energy. We have not tried to study this effect numerically because it requires many ${\bf p}$-modes around the original one, but we have proven energy conservation analytically.
The usual EOMs correctly account for flavor evolution, meaning the trace-free part of the density matrices often described by polarization vectors. On the other hand, the particle number in a given ${\bf p}$-mode is conserved, corresponding to the trace part of the EOM, where the left-hand side of Eq. (1) is a continuity equation. Therefore, the trace part of the EOM must be expanded to higher order in the gradients to capture nontrivial evolution. The reshuffling of neutrinos among ${\bf p}$-modes is a small effect relative to the large kinetic energies, yet precisely absorbs the missing refractive energy. Unless one wants to follow small modifications of the neutrino energy distribution, Eq. (1) remains the one to be studied.
We note that Eq. (1) conserves entropy and including the new gradient terms, this is also the case, i.e., entropy is conserved order by order in the gradient expansion (see Supplemental Material). On the practical level, the conservation of entropy, but not of energy, could be used to test numerical stability.
For the moment, our finding is primarily of conceptual interest: at least superficially, the roles of energy and entropy are somewhat exchanged relative to a naive expectation. This finding may shed new light on a recent proposal that the final outcome of fast conversions may be some sort of thermalized state Johns:2023jjt . The quasi-steady state generically observed in numerical simulations of fast conversions Padilla-Gay:2020uxa ; Zaizen:2023ihz ; Xiong:2023vcm ; Abbar:2023ltx ; Cornelius:2023eop ; Urquilla:2024bvf is only determined by the oscillatory part of the density matrix, yet its dynamics does not admit a separately conserved energy. Thus, it cannot separately thermalize, since it exchanges energy with the kinetic energy of neutrinos.
Acknowledgments.—We acknowledge useful discussion with Basudeb Dasgupta, Lucas Johns, Shashank Shalgar, and Irene Tamborra. DFGF is supported by the Villum Fonden under Project No. 29388 and the European Union’s Horizon 2020 Research and Innovation Program under the Marie Skłodowska-Curie Grant Agreement No. 847523 “INTERACTIONS.” GGR acknowledges partial support by the German Research Foundation (DFG) through the Collaborative Research Centre “Neutrinos and Dark Matter in Astro- and Particle Physics (NDM),” Grant SFB-1258-283604770, and under Germany’s Excellence Strategy through the Cluster of Excellence ORIGINS EXC-2094-390783311. GS acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306.
Supplemental Material for the Letter
Inhomogeneous Kinetic Equation for Mixed Neutrinos: Tracing the Missing Energy
In this Supplemental Material, we provide further details on the derivation of the equation of motion (EOM) discussed in the main text.
Appendix A A. Detailed derivation of the equations of motion
For the reader’s convenience, we reproduce some of the equations from the main text. The starting point of the derivation is the many-body Hamiltonian describing the neutrino kinetic energy and interactions. This is written $\mathcal{H}=\mathcal{H}_{0}+\mathcal{U}$, where $\mathcal{H}_{0}=\sum_{\alpha,{\bf p}}\epsilon_{\bf p}a^{\dagger}_{\alpha,{\bf p}}a_{\alpha,{\bf p}}$ is the neutrino kinetic energy operator, and
$\mathcal{U}=\frac{\sqrt{2}G_{\rm F}}{8}\sum_{\left\{{\bf p}\right\},\alpha,\beta}a^{\dagger}_{\alpha,{\bf p}_{1}}a_{\alpha,{\bf p}_{2}}a^{\dagger}_{\beta,{\bf p}_{3}}a_{\beta,{\bf p}_{4}}\overline{u}_{{\bf p}_{1}}\gamma^{\mu}u_{{\bf p}_{2}}\overline{u}_{{\bf p}_{3}}\gamma_{\mu}u_{{\bf p}_{4}}.$ (S1)
We will henceforth use the notation $({\bf p}_{1},{\bf p}_{2})({\bf p}_{3},{\bf p}_{4})=\overline{u}_{{\bf p}_{1}}\gamma^{\mu}u_{{\bf p}_{2}}\overline{u}_{{\bf p}_{3}}\gamma_{\mu}u_{{\bf p}_{4}}$.
We want to determine the time evolution, under this Hamiltonian, of the density operator
$\mathcal{D}_{\alpha\beta}({\bf p},{\bf r})=\sum_{\bf k}a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a_{\alpha,{\bf p}-{\bf k}/2}e^{-i{\bf k}\cdot{\bf r}},$ (S2)
or more conveniently of its spatial Fourier transform
$\tilde{\mathcal{D}}_{\alpha\beta}({\bf p},{\bf k})=a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a_{\alpha,{\bf p}-{\bf k}/2}.$ (S3)
Following the notation in the main text, we denote the expectation value over the quantum state $\langle\mathcal{D}_{\alpha\beta}({\bf p},{\bf r})\rangle=\rho_{\alpha\beta}({\bf p},{\bf r})$ and analogously for the Fourier transform $\langle\tilde{\mathcal{D}}_{\alpha\beta}({\bf p},{\bf k})\rangle=\tilde{\rho}_{\alpha\beta}({\bf p},{\bf k})$.
We start by obtaining the commutator of the non-interacting Hamiltonian $\mathcal{H}_{0}$ with this operator; using the standard commutation rules we easily find
$[\mathcal{H}_{0},\tilde{\mathcal{D}}_{\alpha\beta}({\bf k},{\bf p})]=\left(\epsilon_{{\bf p}+{\bf k}/2}-\epsilon_{{\bf p}-{\bf k}/2}\right)\tilde{\mathcal{D}}_{\alpha\beta}({\bf k},{\bf p}).$ (S4)
Expanding to first order in ${\bf k}$, and using ${\bm{\nabla}}_{\bf p}\epsilon_{\bf p}={\bf v}$, we find
$\left(\frac{\partial\tilde{\mathcal{D}}_{\alpha\beta}}{\partial t}\right)_{0}=i[\mathcal{H}_{0},\tilde{\mathcal{D}}_{\alpha\beta}]=i{\bf v}\cdot{\bf k}\,\tilde{\mathcal{D}}_{\alpha\beta},$ (S5)
which Fourier transformed leads to the usual advective term quoted in the main text.
Let us now consider the commutator of the interaction Hamiltonian $\mathcal{U}$, which gives
$[\mathcal{U},\tilde{\mathcal{D}}_{\alpha\beta}]$ $=$ $\frac{\sqrt{2}G_{\rm F}}{8}\sum_{\left\{{\bf p}\right\},\gamma}({\bf p}_{1},{\bf p}_{2})({\bf p}_{3},{\bf p}_{4})$ (S6)
×$\kern 50.00008pt{}\times\left[a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\gamma,{\bf p}_{2}}a^{\dagger}_{\beta,{\bf p}_{3}}a_{\alpha,{\bf p}-{\bf k}/2}\delta_{{\bf p}_{4},{\bf p}+{\bf k}/2}+a^{\dagger}_{\beta,{\bf p}_{1}}a^{\dagger}_{\gamma,{\bf p}_{3}}a_{\gamma,{\bf p}_{4}}a_{\alpha,{\bf p}-{\bf k}/2}\delta_{{\bf p}_{2},{\bf p}+{\bf k}/2}\right.$
$\kern 50.00008pt\left.\kern 10.00002pt{}-a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\gamma,{\bf p}_{2}}a_{\alpha,{\bf p}_{4}}\delta_{{\bf p}_{3},{\bf p}-{\bf k}/2}-a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a_{\alpha,{\bf p}_{2}}a^{\dagger}_{\gamma,{\bf p}_{3}}a_{\gamma,{\bf p}_{4}}\delta_{{\bf p}_{1},{\bf p}-{\bf k}/2}\right],$
where $\delta_{{\bf p}_{1},{\bf p}_{2}}$ is a Dirac delta forcing the equality ${\bf p}_{1}={\bf p}_{2}$. We can now take the expectation value of this operator over the mean-field state; the corresponding terms can be grouped in pairs which are identical after the exchange ${\bf p}_{1}\to{\bf p}_{3}$, ${\bf p}_{2}\to{\bf p}_{4}$, so we finally obtain
$\langle[\mathcal{U},\tilde{\mathcal{D}}_{\alpha\beta}]\rangle$ $=$ $\frac{\sqrt{2}G_{\rm F}}{4}\sum_{\left\{{\bf p}\right\},\gamma}({\bf p}_{1},{\bf p}_{2})({\bf p}_{3},{\bf p}_{4})$ (S7)
×$\kern 50.00008pt{}\times\left[(\langle a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\gamma,{\bf p}_{2}}\rangle\langle a^{\dagger}_{\beta,{\bf p}_{3}}a_{\alpha,{\bf p}-{\bf k}/2}\rangle\right.-\langle a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\alpha,{\bf p}-{\bf k}/2}\rangle\langle a^{\dagger}_{\beta,{\bf p}_{3}}a_{\gamma,{\bf p}_{2}}\rangle)\delta_{{\bf p}_{4},{\bf p}+{\bf k}/2}$
$\kern 50.00008pt\left.\kern 10.00002pt{}-(\langle a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\gamma,{\bf p}_{2}}\rangle\langle a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a_{\alpha,{\bf p}_{4}}\rangle-\langle a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\alpha,{\bf p}_{4}}\rangle\langle a^{\dagger}_{\beta,{\bf p}+{\bf k}/2}a_{\beta,{\bf p}_{4}}\rangle)\delta_{{\bf p}_{3},{\bf p}-{\bf k}/2}\right].$
In each term, it will prove convenient to reparametrize the momenta in such a way as to separate the large momenta (associated with the neutrino energies) from the small momenta (associated with the spatial gradients of the density matrix). In particular, in terms of the form $\langle a^{\dagger}_{\beta,{\bf p}_{1}}a_{\alpha,{\bf p}_{2}}\rangle$, because of the slowly-varying nature of the density matrix, we find a non-vanishing result only if ${\bf p}_{1}$ and ${\bf p}_{2}$ differ by terms of order $\sim\ell^{-1}$. Thus, for example, it proves convenient to reparameterize the first term $\langle a^{\dagger}_{\gamma,{\bf p}_{1}}a_{\gamma,{\bf p}_{2}}\rangle\langle a^{\dagger}_{\alpha,{\bf p}_{3}}a_{\beta,{\bf p}-{\bf k}/2}\rangle$ choosing ${\bf p}_{1}={\bf p}^{\prime}+({\bf k}-{\bf q})/4$, ${\bf p}_{2}={\bf p}^{\prime}-({\bf k}-{\bf q})/4$, ${\bf p}_{3}={\bf p}+{\bf q}/2$, ${\bf p}_{4}={\bf p}+{\bf k}/2$, which makes it clear that ${\bf p}$ and ${\bf p}^{\prime}$ are large momenta, while ${\bf q}$ and ${\bf k}$ are small. Adopting a similar parameterization for all the terms we find
$\langle[\mathcal{U},\tilde{\mathcal{D}}_{\alpha\beta}]\rangle$ $=$ $\frac{\sqrt{2}G_{\rm F}}{4}\sum_{{\bf p}^{\prime},{\bf q},\gamma}\left[\tilde{\rho}_{\alpha\beta}\left({\bf p}+\frac{{\bf q}-{\bf k}}{4},\frac{{\bf q}+{\bf k}}{2}\right)\tilde{\rho}_{\gamma\gamma}\left({\bf p}^{\prime},\frac{{\bf k}-{\bf q}}{2}\right)\right.\left({\bf p}+\frac{{\bf q}}{2},{\bf p}+\frac{{\bf k}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf k}-{\bf q}}{4},{\bf p}^{\prime}+\frac{{\bf q}-{\bf k}}{4}\right)$ (S8)
$\kern 20.00003pt{}-\tilde{\rho}_{\alpha\gamma}\left({\bf p}+\frac{{\bf q}-{\bf k}}{4},\frac{{\bf q}+{\bf k}}{2}\right)\tilde{\rho}_{\gamma\beta}\left({\bf p}^{\prime},\frac{{\bf k}-{\bf q}}{2}\right)\left({\bf p}+\frac{{\bf q}}{2},{\bf p}^{\prime}+\frac{{\bf q}-{\bf k}}{4}\right)\left({\bf p}^{\prime}+\frac{{\bf k}-{\bf q}}{4},{\bf p}+\frac{{\bf k}}{2}\right)$
$\kern 20.00003pt{}-\tilde{\rho}_{\alpha\beta}\left({\bf p}+\frac{{\bf k}-{\bf q}}{4},\frac{{\bf k}+{\bf q}}{2}\right)\tilde{\rho}_{\gamma\gamma}\left({\bf p}^{\prime},\frac{{\bf k}-{\bf q}}{2}\right)\left({\bf p}-\frac{{\bf k}}{2},{\bf p}-\frac{{\bf q}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf k}-{\bf q}}{4},{\bf p}^{\prime}+\frac{{\bf q}-{\bf k}}{4}\right)$
$\kern 20.00003pt{}+\tilde{\rho}_{\alpha\gamma}\left({\bf p}^{\prime},\frac{{\bf k}-{\bf q}}{2}\right)\tilde{\rho}_{\gamma\beta}\left({\bf p}+\frac{{\bf k}-{\bf q}}{4},\frac{{\bf k}+{\bf q}}{2}\right)\left.\left({\bf p}-\frac{{\bf k}}{2},{\bf p}^{\prime}+\frac{{\bf q}-{\bf k}}{4}\right)\left({\bf p}^{\prime}+\frac{{\bf k}-{\bf q}}{4},{\bf p}-\frac{{\bf q}}{2}\right)\right].$
We can now take the inverse Fourier transform by multiplying by $e^{-i{\bf k}\cdot{\bf r}_{1}}$ and summing over all ${\bf k}$. We perform the full manipulations only on the first of the four terms; the other ones are treated analogously. Introducing the notation ${\bf a}={\bf k}-{\bf q}$, $\mathbf{b}={\bf k}+{\bf q}$, we can write this term as
$\sum_{\frac{{\bf a}}{2},\frac{\mathbf{b}}{2},{\bf p}^{\prime},\gamma}e^{i\frac{\mathbf{b}}{2}\cdot({\bf r}_{2}-{\bf r}_{1})+i\frac{{\bf a}}{2}\cdot({\bf r}_{3}-{\bf r}_{1})}\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r}_{3})\rho_{\alpha\beta}\left({\bf p}-\frac{{\bf a}}{4},{\bf r}_{2}\right)\left({\bf p}+\frac{{\bf q}}{2},{\bf p}+\frac{{\bf k}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf a}}{4},{\bf p}^{\prime}-\frac{{\bf a}}{4}\right).$
In this expression, we must now perform the expansion for $|{\bf k},{\bf q}|\ll|{\bf p},{\bf p}^{\prime}|$. For the spinor part, this requires using the explicit form of the spinors expanded to first order in the momenta ${\bf a}$ and $\mathbf{b}$. The result depends on the mutual orientation of the vectors ${\bf p}$ and ${\bf p}^{\prime}$. Thus, we introduce a set of three orthonormal directions
${\bf e}_{y}=\frac{{\bf p}\times{\bf p}^{\prime}}{pp^{\prime}\sin\theta_{{\bf p}{\bf p}^{\prime}}},\quad{\bf e}_{x}=\frac{{\bf e}_{y}\times{\bf p}}{p},\quad{\bf e}_{z}=\frac{{\bf p}}{p}.$ (S9)
We can now use the result that
$\left({\bf p}+\frac{{\bf q}}{2},{\bf p}+\frac{{\bf k}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf a}}{4},{\bf p}^{\prime}-\frac{{\bf a}}{4}\right)\simeq 4(1-\cos\theta_{{\bf p}{\bf p}^{\prime}})-\frac{\sin\theta_{{\bf p}{\bf p}^{\prime}}[b_{x}p^{\prime}+ia_{y}(p+p^{\prime})]}{pp^{\prime}}.$ (S10)
This expression is obtained by a direct use of the corresponding spinors. The other matrix element is
$\left({\bf p}+\frac{{\bf q}}{2},{\bf p}^{\prime}-\frac{{\bf a}}{4}\right)\left({\bf p}^{\prime}+\frac{{\bf a}}{4},{\bf p}+\frac{{\bf k}}{2}\right)=-\left({\bf p}+\frac{{\bf q}}{2},{\bf p}+\frac{{\bf k}}{2}\right)\left({\bf p}^{\prime}+\frac{{\bf a}}{4},{\bf p}^{\prime}-\frac{{\bf a}}{4}\right)$ (S11)
by virtue of Fierz identities.
The expansion in powers of ${\bf k}$ and ${\bf q}$ involves three terms. The zeroth-order term is then simply
$4\sum_{{\bf p}^{\prime},\gamma}(1-\cos\theta_{{\bf p}{\bf p}^{\prime}})\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r})\rho_{\alpha\beta}({\bf p},{\bf r}),$ (S12)
where we call ${\bf r}_{1}={\bf r}$ since only one spatial variable survives in the final result. Next we have a first-order term coming from the gradient expansion $\rho_{\alpha\beta}({\bf p}-{\bf a}/4,{\bf r}_{2})\simeq\rho_{\alpha\beta}({\bf p},{\bf r}_{2})-{\bm{\nabla}}_{\bf p}\rho_{\alpha\beta}\cdot{\bf a}/4$, leading to
$-2i\sum_{{\bf p}^{\prime},\gamma}(1-\cos\theta_{{\bf p}{\bf p}^{\prime}}){\bm{\nabla}}_{\bf r}\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r})\cdot{\bm{\nabla}}_{\bf p}\rho_{\alpha\beta}({\bf p},{\bf r}).$ (S13)
Finally, from the expansion of the spinor part we obtain
$-\sum_{{\bf p}^{\prime},\gamma}\frac{\sin\theta_{{\bf p}{\bf p}^{\prime}}}{pp^{\prime}}\Bigl{[}2ip^{\prime}\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r})\,{\bf e}_{x}\cdot{\bm{\nabla}}_{\bf r}\rho_{\alpha\beta}({\bf p},{\bf r})-2(p+p^{\prime})\rho_{\alpha\beta}({\bf p},{\bf r})\,{\bf e}_{y}\cdot{\bm{\nabla}}_{\bf r}\rho_{\gamma\gamma}({\bf p}^{\prime},{\bf r})\Bigr{]}.$ (S14)
Performing analogous manipulations on all four terms of Eq. (S7) we finally reach the form
$\frac{\partial\rho_{{\bf p},{\bf r}}}{\partial t}+{\bf v}\cdot{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}=i\sqrt{2}G_{\rm F}\sum_{{\bf p}^{\prime}}(1-\cos\theta_{{\bf p}{\bf p}^{\prime}})[\rho_{{\bf p},{\bf r}},\rho_{{{\bf p}^{\prime},{\bf r}}}]+\left(\frac{\partial\rho_{{\bf p},{\bf r}}}{\partial t}\right)_{\mathrm{der}}+\left(\frac{\partial\rho_{{\bf p},{\bf r}}}{\partial t}\right)_{\mathrm{spinor}},$ (S15)
where we introduce the compact notation $\rho_{{\bf p},{\bf r}}=\rho({\bf p},{\bf r})$ as in the main text. The terms $(\partial\rho_{{\bf p},{\bf r}}/\partial t)_{\mathrm{der}}$ originate from the derivatives of the density matrices with respect to momenta, analogously to Eq. (S13), and can be written as
$\left(\frac{\partial\rho_{{\bf p},{\bf r}}}{\partial t}\right)_{\mathrm{der}}=\frac{\sqrt{2}G_{\rm F}}{2}\sum_{{\bf p}^{\prime}}(1-\cos\theta_{{\bf p}{\bf p}^{\prime}})\left(\Bigl{\{}{\bm{\nabla}}_{\bf r}\rho_{{\bf p}^{\prime},{\bf r}},{\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\Bigr{\}}+2\mathrm{Tr}({\bm{\nabla}}_{\bf r}\rho_{{\bf p}^{\prime},{\bf r}})\cdot{\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\right).$ (S16)
The expansion of the spinors leads to a term of the form
$\left(\frac{\partial\rho_{{\bf p},{\bf r}}}{\partial t}\right)_{\mathrm{spinor}}$ $=$ $\frac{\sqrt{2}G_{\rm F}}{2}\sum_{{\bf p}^{\prime}}\frac{\sin\theta_{{\bf p}{\bf p}^{\prime}}{\bf e}_{x}}{p}\cdot\left(\Bigl{\{}\rho_{{\bf p}^{\prime},{\bf r}},{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}\Bigr{\}}+2\mathrm{Tr}(\rho_{{\bf p}^{\prime},{\bf r}}){\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}\right)$ (S17)
${}+\frac{i\sqrt{2}G_{\rm F}}{2}\sum_{{\bf p}^{\prime}}\sin\theta_{{\bf p}{\bf p}^{\prime}}\left(\frac{1}{p}+\frac{1}{p^{\prime}}\right){\bf e}_{y}\cdot\Bigl{[}\rho_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf r}\rho_{{\bf p}^{\prime},{\bf r}}\Bigr{]}.$
Using now the identity
${\bm{\nabla}}_{\bf p}\cos\theta_{{\bf p}{\bf p}^{\prime}}=\frac{\sin\theta_{{\bf p}{\bf p}^{\prime}}}{p}\,{\bf e}_{x},$ (S18)
and repeatedly integrating by parts, we are led to the equations finally reported in the main text.
Appendix B B. Entropy Conservation
Besides conserving energy, the newly introduced EOMs have also the property of conserving the entropy of the neutrino gas, defined by
$S_{\nu}=\int d^{3}{\bf r}\sum_{{\bf p}}\mathrm{Tr}\left(s_{{\bf p},{\bf r}}\right),$ (S19)
where
$s_{{\bf p},{\bf r}}=-\Bigl{[}\rho_{{\bf p},{\bf r}}\log\rho_{{\bf p},{\bf r}}+(\mathbb{1}-\rho_{{\bf p},{\bf r}})\log(\mathbb{1}-\rho_{{\bf p},{\bf r}})\Bigr{]},$ (S20)
so that the time derivative can be written
$\frac{dS_{\nu}}{dt}=\int d^{3}{\bf r}\sum_{{\bf p}}\mathrm{Tr}\left[\frac{\partial\rho_{{\bf p},{\bf r}}}{\partial t}\log\left(\frac{\mathbb{1}-\rho_{{\bf p},{\bf r}}}{\rho_{{\bf p},{\bf r}}}\right)\right].$ (S21)
For the standard EOMs, without the new terms, entropy conservation is automatically enforced. We now show that the new terms from the gradient expansion also maintain this conservation law.
To compute the time derivative of the entropy, it is convenient to rewrite the EOMs reported in the main text as
$\partial_{t}\rho_{{\bf p},{\bf r}}+{\bf v}\cdot{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}=i\Bigl{[}\rho_{{\bf p},{\bf r}},{\sf\Omega}_{{\bf p},{\bf r}}\Bigr{]}+\frac{1}{2}\left\{{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\right\}-\frac{1}{2}\left\{{\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf r}\rho_{{\bf p},{\bf r}}\right\}.$ (S22)
We can now replace the time derivative Eq. (S22) in the rate of change of the entropy Eq. (S21); the advection term of course does not lead to any change in entropy. The commutator term also leads to a vanishing entropy change, as follows from the identity $\mathrm{Tr}\left([{\sf\Omega}_{{\bf p},{\bf r}},\rho_{{\bf p},{\bf r}}]\,{\sf F}(\rho_{{\bf p},{\bf r}})\right)=0$, where the matrix ${\sf F}$ is a generic function. Finally, from the anticommutator terms we find
$\frac{dS_{\nu}}{dt}=\frac{1}{2}\int d^{3}{\bf r}\sum_{{\bf p}}\mathrm{Tr}\left[\left\{{\bm{\nabla}}_{\bf r}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf p}s_{{\bf p},{\bf r}}\right\}-\left\{{\bm{\nabla}}_{\bf p}{\sf\Omega}^{(0)}_{{\bf p},{\bf r}},{\bm{\nabla}}_{\bf r}s_{{\bf p},{\bf r}}\right\}\right],$ (S23)
where we have recognized the total derivative
${\bm{\nabla}}_{\bf p}\rho_{{\bf p},{\bf r}}\log\left(\frac{\mathbb{1}-\rho_{{\bf p},{\bf r}}}{\rho_{{\bf p},{\bf r}}}\right)={\bm{\nabla}}_{\bf p}s_{{\bf p},{\bf r}}.$ (S24)
Even though the operators $\rho_{{\bf p},{\bf r}}$ and $h_{{\bf p},{\bf r}}$ do not commute, this manipulation can be performed for all terms within the trace. Finally, by integrating by parts both terms, we can easily conclude that the rate of change of the entropy vanishes. Therefore, we recover the physically intuitive conclusion that, in the absence of incoherent collisions, which could only appear to second order in $G_{\rm F}$, neutrino entropy is conserved.